Use the confidence interval to find the estimated margin of error. Then find the sample mean.

A store manager reports a confidence interval of

left parenthesis 42.1 comma 78.9 right parenthesis(42.1,78.9)

when estimating the mean price​ (in dollars) for the population of textbooks.

The estimated margin of error is

nothing.
Answer #1

Solution:

Confidence interval = Sample mean ± Margin of error

Confidence interval = Xbar ± E

Lower limit = Xbar – E = 42.1

Upper limit = Xbar + E = 78.9

Xbar = (Lower limit + Upper limit) / 2

Xbar = (42.1 + 78.9) / 2

Xbar = 60.5

Margin of error = E = (Upper limit – lower limit) / 2

Margin of error = (78.9 - 42.1)/2 = 18.4

Margin of error = 18.4

The estimated margin of error is 18.4.

The sample mean is 60.5.